OpenStudy (anonymous):

what is the correct classification of this system of equations? −2x+y=7ˏ3y=6x+4

OpenStudy (anonymous):

yes thank you

OpenStudy (texaschic101):

wait...I was wrong

OpenStudy (texaschic101):

3y = 6x + 4 ---> -6x + 3y = 4 -6x + 3y = 4 -2x + y = 7 --->(-3) ----------------- -6x + 3y = 4 6x - 3y = -21 (result of multiplying by -3) ----------------add 0 = - 17 There is NO SOLUTION to this system

OpenStudy (anonymous):

f(x)=(2x+3)(x-2) and g(x)=2x^2-16. find g(x)-f(x)